What is the wavelenth of light emitted when the electron in ahydrogen atom undergoes transition from an energy level with N=4 to an energy level with N=2?

Grade:11

1 Answers

Arun
6 years ago

According to Balmer formula

Wave no = RH [1/n12 – 1/n22]

Here n1 = 2 ,       n2 = 4,     RH = 109678

Putting these values in the equation we get

Wave number =109678(1/2- ¼2) = 109678 x 3/16

Also λ = 1/ wave number

Therefore λ = 16 / 109678 x 3 = 486 nm